# How to Learn a Star: Binary Classification with Starshaped Polyhedral Sets

### v1

@@ Appendix B Proofs
We show this statement by giving an example of a data-set with a disconnected level set $L(\operatorname{err},0)$. For this, we continue with Section A.3, a configuration of $3$ data points in $\mathbb{R}^{2}$. The $2$-dimensional Coxeter fan of type B has $8$ rays and $8$ maximal cones. Thus, the parameter space >$\mathbb{R}^{8}_{>0}\times\mathbb{R}^{2}$ is subdivided into cells of the form $\bigcap_{C\in\Delta}\ \bigcap_{i\in\{1,2,3\}}\mathcal{S}^{b(C,i)}(C,\mathbf{x}^{(i)})$, many of these are empty or lower dimensional. The extended 0/1-loss attains the value 00 on $16$ maximal cells. To describe them, we use the indexing of the rays as depicted in Figure 3(b), and denote $C_{i}=\operatorname{cone}(\mathbf{v}_{i},\mathbf{v}_{i+1})$ for $i=1,\dots,7$, $C_{8}=\operatorname{cone}(\mathbf{v}_{8},\mathbf{v}_{1})$. One example of a valid configuration is a tuple $(\mathbf{a},\mathbf{t})$ such that $\mathbf{x}^{(1)}\in(\operatorname{star}(\mathbf{a})+\mathbf{t})\cap(C_{4}+\mathbf{t}),\ \mathbf{x}^{(2)}\in(\mathbb{R}^{2}\setminus(\operatorname{star}(\mathbf{a})+\mathbf{t}))\cap(C_{3}+\mathbf{t})$ and $\mathbf{x}^{(3)}\in(\operatorname{star}(\mathbf{a})+\mathbf{t})\cap(C_{2}+\mathbf{t})$, as depicted in Figure 7(a). The set of $(\mathbf{a},\mathbf{t})$ satisfying these conditions is $\mathcal{S}^{0}(C_{4},\mathbf{x}^{(1)})\cap\mathcal{S}^{1}(C_{3},\mathbf{x}^{(2)})\cap\mathcal{S}^{0}(C_{2},\mathbf{x}^{(3)})$. In total, the set of perfect classifiers, i.e., the $0^{\text{th}}$ level set is the union of the following $16$ nonempty cells:
@@
It can be checked that the union of the sets in each of the above columns is path connected. However, there is no path from a point in the left column to a point in the right column. If these sets were path connected, then there was a path from the configuration depicted in Figure 7(a) to the configuration in Figure 7(b) by a continuous translation of the star and by shifting the points $\tfrac{1}{a_{i}}\mathbf{v}_{i}$ along the rays continuously. But since $\mathbf{x}^{(2)}$ lies in the convex hull of $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(3)}$ and the line segment through any of these point is parallel to the rays $\mathbf{v}_{2},\mathbf{v}_{6}$, any such continuous transformation necessarily increases the number of mistakes at a certain point. ∎
B.14 Proof of Theorem 4.5
@@
By Kupavskii [2020], the set of all simplices has VC dimension $O(d^{2}\log_{2}(d))$. Any starshaped classifier $c_{\mathbf{a},\mathbf{t}}$ is a union of $k$ simplices. Thus, the result follows from Blumer et al. [1989]. ∎
Marie-Charlotte Brandenburg
Ruhr-Universität Bochum
Universitätsstraße 150, 44801 Bochum, Germany
marie-charlotte.brandenburg@rub.de
Katharina Jochemko
KTH Royal Institute of Technology
SE-100 44 Stockholm
jochemko@kth.se

### v2

@@ Appendix C Proofs
We show this statement by giving an example of a data-set with a disconnected level set $L(\operatorname{err},0)$. For this, we continue with Section˜A.3, a configuration of $3$ data points in $\mathbb{R}^{2}$. The $2$-dimensional Coxeter fan of type B has $8$ rays and $8$ maximal cones. Thus, the parameter space $\mathbb{R}^{8}_{>0}\times\mathbb{R}^{2}$ is subdivided into cells of the form $\bigcap_{C\in\Delta}\ \bigcap_{i\in\{1,2,3\}}\mathcal{S}^{b(C,i)}(C,\mathbf{x}^{(i)})$, many of these are empty or lower dimensional. The extended 0/1-loss attains the value 0 on $16$ maximal cells. To describe them, we use the indexing of the rays as depicted in Figure˜4(b), and denote $C_{i}=\operatorname{cone}(\mathbf{v}_{i},\mathbf{v}_{i+1})$ for $i=1,\dots,7$, $C_{8}=\operatorname{cone}(\mathbf{v}_{8},\mathbf{v}_{1})$. One example of a valid configuration is a tuple $(\mathbf{a},\mathbf{t})$ such that $\mathbf{x}^{(1)}\in(\operatorname{star}(\mathbf{a})+\mathbf{t})\cap(C_{4}+\mathbf{t}),\ \mathbf{x}^{(2)}\in(\mathbb{R}^{2}\setminus(\operatorname{star}(\mathbf{a})+\mathbf{t}))\cap(C_{3}+\mathbf{t})$ and $\mathbf{x}^{(3)}\in(\operatorname{star}(\mathbf{a})+\mathbf{t})\cap(C_{2}+\mathbf{t})$, as depicted in Figure˜8(a). The set of $(\mathbf{a},\mathbf{t})$ satisfying these conditions is $\mathcal{S}^{0}(C_{4},\mathbf{x}^{(1)})\cap\mathcal{S}^{1}(C_{3},\mathbf{x}^{(2)})\cap\mathcal{S}^{0}(C_{2},\mathbf{x}^{(3)})$. In total, the set of perfect classifiers, i.e., the $0^{\text{th}}$ level set is the union of the following $16$ nonempty cells:
@@
It can be checked that the union of the sets in each of the above columns is path connected. However, there is no path from a point in the left column to a point in the right column. If these sets were path connected, then there was a path from the configuration depicted in Figure˜8(a) to the configuration in Figure˜8(b) by a continuous translation of the star and by shifting the points $\tfrac{1}{a_{i}}\mathbf{v}_{i}$ along the rays continuously. But since $\mathbf{x}^{(2)}$ lies in the convex hull of $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(3)}$ and the line segment through any of these point is parallel to the rays $\mathbf{v}_{2},\mathbf{v}_{6}$, any such continuous transformation necessarily increases the number of mistakes at a certain point. ∎
C.14 Proof of Theorem˜4.9
@@
By Kupavskii_2020, the set of all simplices has VC dimension $O(d^{2}\log_{2}(d))$. Any starshaped classifier $c_{\mathbf{a},\mathbf{t}}$ is a union of $k$ simplices. Thus, the result follows from blumer1989learnability. ∎